Grade 11IIT JEE Entrance Exam

2logx(base 3)-4logx(base 27)
pls answer quickly with detailed steps pls answer as soon as possible

Sathvik Mp Rao
9 Years agoGrade 11

1 Answer

Manish Tiwari
9 Years ago
Your question is-2logx(base3) - 4logx(base 27)=2logx(base3) - 4/3logx(base 3)=(2-4/3)logx(base3)=2/3logx(base 3)Hope it helps!!